What is the mass of water if it contains 6.022 x 1025 atoms of oxygen


1
Expert's answer
2022-02-17T03:50:02-0500

Molecular formula for water is "H_2O"

Mass of oxygen =16g which is equivalent to "6.022\times10^{25}atoms"

Atoms of hydrogen "=2/16\times6.022\times 10^{25}=7.5275\times 10^{24} atoms"

Total atoms in water "=6.022\times 10^{25} +7.5275\times 10^{24}=6.77475\times 10^{25}atoms"

Avogadro's number = "6.022\times 10^{23}"

Mass of water "=\frac{6.77475\times 10^{25}}{6.022\times 10^{23}}=112.5g"